A quadratic equation has two roots: 3/4 and -5. Find a quadratic equation where the coefficient of the x^2 term is 1 and find a second equation that has only integers as coefficients.
Answer by lwsshak3(11628)

A quadratic equation has two roots: 3/4 and -5. Find a quadratic equation where the coefficient of the x^2 term is 1 and find a second equation that has only integers as coefficients.
**
(x-3/4)(x+5)=0
expand
x^2-3x/4+5x-15/4=0
x^2+17x/4-15/4=0 (coefficient of x^2=1)
multiply by LCD:4
4x^2+17x-15=0 (coefficients are all integers)